Peter cycles at 4 mph uphill, and 12 mph downhill. He rides from his house at the bottom of a hill to his grandma’s house at the top of the hill, and back home. His average speed for the whole journey is:
Answer by Alan3354(69443)

Peter cycles at 4 mph uphill, and 12 mph downhill. He rides from his house at the bottom of a hill to his grandma’s house at the top of the hill, and back home. His average speed for the whole journey is:
---------
Avg = 2*4*12/(4+12) = 96/16
Avg = 6 mi/hr